How many grams of silver that is 60% pure must be mixed together with silver that is 35% pure to obtain a mixture of 90 grams of silver that is 45% pure?
Answer by rfer(16322)

60%=x=36 gr
35%=90-x=54 gr
.6x+.35(90-x)=.45*90
.6x+31.5-.35x=40.5
.25x=9
x=36